how could i find "x" in this problem? 9/(x-3) - (x-4)/(x-3) = 1/4
Answer by stanbon(75887)

9/(x-3) - (x-4)/(x-3) = 1/4
Multiply through by 4(x-3) to get;
4*9 - 4(x-4) = x-3
36 -4x+16=x-3
5x=55
x=11
Cheers,
Stan H.